X+84 = x - 29
_____. _
70. 61. Solving equal fractions with un common denominators

Found 2 solutions by Boreal, solver91311:
Answer by Boreal(15235)

(x+84)/70= (x-29)/61
Cross-multiply
(x+84)*61=(x-29)*70
61x+5124=70x-2030
7154=9x
x=794 8/9=794.89
======
794.888+84=878.888.
Divide by 70=12.555
765.89/61=12.555
x=794 8/9

Answer by solver91311(24713)



If you are trying to write , then

First, render your equation in plain text like this:

(x + 84)/(70) = (x - 29)/(61).

That we understand; the mess you made, not so much. One more thing: X and x are two different things; be consistent with your capitalization.

Second, this is a proportion, so cross multiply: Numerator on the left times denominator on the right is equal to denominator on the left times numerator on the right. That gives you a linear equation in that can be solved by ordinary means.

See Formatting Math as Text for more information about how to communicate math expressions in plain text that everyone understands.

John

My calculator said it, I believe it, that settles it